Quantum.  In an experiment, a photon's frequency is measured to be 4 x 1016 Hz.  What is the energy of the photon, in units of electron-Volts?
Answer #1

Givein in the question,

Frequency of the photon, n = 4 × 1016 Hz

Now, energy of a photon havin frequency n Hz is,

E = h × n

Hare, h= planck's constant

= 6.626 × 10-34 J.s

Thus energy in electron volt,

E = (h×n)/e = ( 6.626 × 10-34 × 4 × 1016)/ (1.6 × 10-19 ) = 165.65 electronvolt.

Answer, E = 165.65 eV.